# DeepEdit: Knowledge Editing as Decoding with Constraints

### v2

@@ Abstract
We propose a new perspective of knowledge editing (KE) for large language models (LLMs) that treats it as a constrained decoding problem. We design decoding constraints to regulate LLMs, ensuring coherence between reasoning steps when incorporating new knowledge. To enforce these constraints, we utilize a depth-first search to adaptively substitute new knowledge for the LLMs’ original reasoning steps, greedily seeking the optimal path of multi-hop reasoning with new knowledge. From this vantage, we propose DeepEdit: Depth-first Search-based Decoding for Knowledge Editing. DeepEdit improves the KE of LLMs by enhancing the conciseness, coherence, pertinence, and receptiveness of reasoning with new knowledge. DeepEdit is flexibly applicable to any black-box LLM without requiring access to model parameters or token-wise distributions. In addition to DeepEdit, we propose two new KE benchmarks: MQuAKE-2002 and MQuAKE-hard, which are designed to provide more precise and challenging assessments of KE approaches. Qualitatively, DeepEdit enables LLMs to produce more succinct reasoning outputs in accordance with new knowledge. Quantitatively, it yields significant improvements on multiple KE benchmarks.
@@ 1 Introduction
Knowledge editing (KE) aims to enable large language models (LLMs) to produce outputs in accordance with the knowledge that is new or conflicted with the knowledge memorized by the LLMs. A challenging task to evaluate KE is the multi-hop question answering with new knowledge Zhong et al. (2023). In this task, LLMs should answer the questions corresponding to a chain of facts including new knowledge (see Figure 1(a)). It is essential for LLMs to deductively infer from the new knowledge in the multi-step reasoning in order to provide correct answers.
Figure 1: (a) An instance of multi-hop question answering with new knowledge. (b) The existing method MeLLo relies on the LLM generator to understand new knowledge by itself. (c) Our DeepEdit applies the constraints to the LLMs’ decoding so as to improve reasoning with new knowledge.
Figure 2: Our DeepEdit enforces LLMs to follow the decoding constraints of [Conciseness], [Coherence], [Receptiveness], and [Pertinence] so as to soundly incorporate new knowledge into LLMs’ reasoning.
Recent literature has mainly sought to perform KE through in-context editing Zhong et al. (2023); Cohen et al. (2023); Zheng et al. (2023a). However, in-context editing lacks direct control of the outputs; the new knowledge‘s influences are not traceable; and the performance suffers from the input-conflicting hallucinations Zhang et al. (2023b) (see Figure 1(b)).
Taking a closer look at the ground-truth output in Figure 1(a), we notice that: when reasoning with new knowledge, LLMs should orderly place new knowledge at the appropriate positions in the reasoning chain, i.e., the new knowledge should coherently bridge the adjacent reasoning steps, and avoid the memorized knowledge that is conflicted to new knowledge. This finding raises a natural question:
@@
The central idea of this paper is to view KE as a problem of constrained decoding. We design the decoding constraints, by following which LLMs can soundly incorporate new knowledge in LLMs reasoning. Specifically, we propose the following constraints: [Conciseness] prevents the redundant loops of reasoning; [Coherence] guarantees the coherence of adjacent steps; [Receptiveness] guarantees the LLMs’ awareness of new knowledge; [Pertinence] improves the relevance of reasoning to the input question (see Figure 4). These constraints improve LLMs’ reasoning with new knowledge from different perspectives as a whole.
Much research has proposed constrained decoding methods Meng et al. (2022c); Liu et al. (2023); Geng et al. (2023). However, classical constrained decoding methods usually require access to the token-wise distribution, which is not provided by many black-box LLMs like ChatGPT OpenAI (2022). In our work, to apply our constraints to the black-box LLMs’ decoding, we breakthrough this restriction by changing the granularity of decoding from tokens to the reasoning steps. We design all decoding actions over the step level, which bypasses the requirement of accessing the token-wise distributions.
At every iteration, we take one generated reasoning step and a few retrieved new facts as the step candidates. We verify the step candidates and filter out the ones that do not satisfy our constraints. Since there could be multiple candidates that satisfy all constraints every iteration. We operate the depth-first search that store all the valid step candidates in a stack and greedily search for the optimal path of LLMs’ multi-hop reasoning with new knowledge. We name our method as DeepEdit, the depth-first search-based decoding for knowledge editing.
The advantages of our DeepEdit on KE are three-folds. First, it does not need to retrain the LLMs like the classical model editing methods Zhu et al. (2020), which is complex and expensive. Second, the new knowledge directly contributes to the LLMs’ decoding by acting as the reasoning step candidates. This facilitates more effective reasoning with new knowledge. Third, the influence of new knowledge is monitored with decoding constraints, leading to more controllable and traceable reasoning with new knowledge. Overall, our DeepEdit enhances the LLMs’ reasoning on new knowledge through higher conciseness, coherence, pertinence, and receptiveness.
DeepEdit can be flexibly applied to all the black-box LLMs. We evaluate DeepEdit on the KE benchmarks MQuAKE-3k, MQuAKE-2002, and MQuAKE-hard Zhong et al. (2023); Meng et al. (2022a). The latter two are new benchmarks that we build to provide more precise and challenging assessments of KE methods. Qualitatively, DeepEdit leads to more succinct and faithful reasoning of LLMs that soundly incorporate new knoweldge, as shown in Figure 8. We also observe substantial quantitative improvements in the KE question-answering accuracy achieved by DeepEdit. Overall, DeepEdit improves the KE for the strong LLMs text-davinci-003, gpt-3.5-turbo-instruct, and llama2 by a significant margin OpenAI (2022); Touvron et al. (2023).
@@ 2 Related Work
Figure 3: Comparison of different KE methods. (upper) Model editing was the most popular KE paradigm, which edits the model parameters to update knowledge. (middle) Recently, in-context editing has become popular, which relies on the LLMs’ capability of utilizing new knowledge in the prompts. (lower) We propose DeepEdit for KE, which directly controls the LLMs’ decoding to follow new knowledge.
Knowledge Editing
Recent work observes the superior KE performance of in-context editing. MeLLo Zhong et al. (2023) designs a single prompt to alternately conduct the text generation and new knowledge analysis using the LLM generator. Cohen et al. (2023) propose to append the new knowledge the head of the input prompt. In-context editing lacks direct control of the LLMs’ outputs and how the new knowledge influences the outputs is not traceable. Different from them, we propose the new decoding method DeepEdit for KE, which effectively augments LLMs on KE during the decoding stage. We illustrate a comparison between DeepEdit and existing KE methods in Figure 3.
Figure 4: Comparison between classical constrained decoding (at the token level) (upper) and our decoding at the reasoning step level (lower). Differences: 1) The token candidates is the predefined vocabulary, while the step candidates include a generated step and the retrieved facts. 2) The token candidates are ranked by generators’ outputs, while the step candidates are filtered by the constraints.
Self-refinement of Large Language Models
Some work has explored to use LLMs to revise the output of themselves in dialogue response, sentiment reversal, and mathematical reasoning Madaan et al. (2023); Shridhar et al. (2023b, a); Welleck et al. (2022); Zheng et al. (2023b); Madaan et al. (2023). However, it is difficult to rely on LLMs to accommodate the new knowledge by themselves in the self-refinement, since the new knowledge is generally conflicted with the parametric knowledge memorized by LLMs. Different from them, we do not rely on LLMs to revise the outputs by themselves. We directly control the decoding to soundly incorporate new knowledge with decoding constraints. We modify the LLMs’ outputs using knowledge retrieval and depth-first search to satisfy the constraints. We hope that our work demonstrates more potential to conduct KE for LLMs in the decoding stage.
@@ 3 Methodology
Multi-hop question answering with new knowledge is a challenging task for assessing knowledge editing (KE) Zhong et al. (2023), where LLMs should answer the questions corresponding to a chain of facts including new knowledge. This task assumes a database storing all new facts that LLMs should follow, and only a few facts can influence the ground-truth answer to a specific question.
Our KE idea is to first design the decoding constraints that can regulate LLMs to soundly incorporate new knowledge in reasoning and then enforce LLMs’ decoding to follow the constraints. In this section, we will introduce how we incorporate our idea into the design of a new KE method DeepEdit (Depth-first Search based Decoding for Knowledge Editing) to improve the KE of black-box LLMs.
3.1 Decoding at the Reasoning Step Level
Many classical constrained decoding methods Zhang et al. (2023a); Meng et al. (2022c) use an auxiliary model to adjust the token-wise distribution during decoding. They manipulate the distribution before every token’s generation to satisfy the constraints. Therefore, they require accessibility to the LLMs’ output distribution, which is not feasible for many black-box LLMs.
In this work, we change the granularity of decoding from the token level to the reasoning step level. Our decoding is designed at the step level: instead of taking the tokens predefined in a fixed vocabulary as the candidates, we take the reasoning steps as the candidates. At every iteration, we let LLMs to produce one reasoning step and retrieve a fixed number of new facts in parallel to the reasoning step as the step candidates. The new fact retrieval is based on the nearest neighbor search with the sentence-level embeddings given by BERT. The motivation of this design is that the new facts that are semantically close to the generated facts are more likely to satisfy the constraints.
Taking the single generated step and the new facts as the step candidates, we filter a valid candidate and take it as the new reasoning step. Note that there may exist more than one candidates that satisfy the constraints. We will introduce how we design the decoding constraints for KE in Section 3.2, how we verify the constraints in Section 3.3 and Section 3.4, and how we process the multiple valid step candidates to search for the optimal path of LLMs’ multi-hop reasoning in Section 3.5.
Our decoding design bypasses the requirement of accessing the token-wise distributions, since all the decoding actions are defined over the step level. Our decoding strategy is free to be applied to black-box LLMs in order to produce satisfactory outputs in accordance with new knowledge. The details are introduced next.
Figure 5: (left) At every decoding iteration, we store the reasoning step candidates (including the generated step and the retrieved knowledge) that satisfy the local constraints in a stack. (right) The global verifier determines the feasibility of the whole reasoning process by checking whether the answer is found. If not feasible, we pop the status stack to use another reasoning subprocess that satisfies the local verifier.
3.2 Designing Decoding Constraints for Knowledge Editing
To correctly answer multi-hop questions in accordance with new knowledge Zhong et al. (2023), LLMs need to generate a sequence of reasoning steps to find the answer. We propose the following constraints for the LLMs’ reasoning with KE.
@@ 3.2 Designing Decoding Constraints for Knowledge Editing
    [Conciseness] Every reasoning step should be different from the prior ones.
@@
    [Coherence] Every reasoning step should be relevant to the last reasoning step.
@@
    [Receptiveness] Every reasoning step cannot contradict to any new knowledge.
@@
    [Pertinence] Every reasoning step should be relevant and helpful to finding the answer.
The above constraints are not automatically satisfied by LLMs when reasoning with new knowledge. However, they are essential to helping LLMs to find the correct answers in accordance with new knowledge. Therefore, in our work, we propose a new decoding method to enforce the LLMs to follow the above constraints so as to improve the KE performance. We first analyze the constraints in detail.
Among the above four constraints, [Conciseness] guarantees that the LLMs’ reasoning is not in a loop. [Coherence] guarantees the coherence of reasoning, while [Pertinence] guarantees the relevance of every reasoning step to the question. [Receptiveness] ensures that LLMs can utilize new knowledge for reasoning. We visualize different constraints in Figure 4. Among the constraints, we regard [Conciseness], [Coherence], and [Receptiveness] as local ones: we can verify whether they are satisfied given only partial reasoning steps. The [Pertinence] constraint is regarded as a global constraint: we can verify it based on a complete reasoning process.
By following the above constraints, LLMs can orderly place new knowledge at appropriate positions in the reasoning chain and avoid the memorized knowledge that is conflicted with new knoweledge. Next, we will introduce how we verify the above constraints and control the LLMs’ decoding to satisfy the constraints.
3.3 Verifier for Local Constraints
To verify the [Conciseness] constraint, by default, whenever a new reasoning step is given, we may compare it with all the prior reasoning steps. However, this is no doubt an expensive solution at the cost of $\mathcal{O}(r^{2})$ times of verification given as many as $r$ reasoning steps.
We observe that the repeated reasoning step generally occurs in the adjacent steps. Therefore, to reduce the computational costs, we propose an approximate verification method to verify the constraint of [Conciseness]. Specifically, for every reasoning step, we only compare it with the last reasoning step, so that the whole time of doing the verification is $\mathcal{O}(r)$.
To satisfy the [Coherence] constraint, we need to evaluate whether the current reasoning step is relevant to the last reasoning step. However, it is challenging to directly and precisely evaluate the relevance through LLMs. We use the LLM based verifiers equipped with a hand-crafted prompt to verify whether the adjacent reasoning steps are semantically equivalent or related to verify the constraints of [Conciseness] and [Coherence] respectively.
Verifying the [Receptiveness] constraint, by default, requires determining whether the contradiction exists in the pair of the reasoning step and all the items in the basis of new knowledge. The time complexity is $\mathcal{O}(rk)$, where $k$ is the number of new knowledge items and a large value in practice. Therefore, to reduce the time complexity, we retrieve the relevant knowledge items that are semantically closest to the reasoning step and then assess the contradictions. The idea is that if two sentences are contradictory to each other, they must first be semantically relevant. In this way, we reduce the time complexity from $\mathcal{O}(rk)$ to $\mathcal{O}(r\log(k))$. Also, we design a LLM-based verifier with a hand-crafted prompt to assess whether the retrieved facts are contradictory to the generated reasoning step.
3.4 Verifier for Global Constraints
Whether a reasoning step is helpful in finding the final answer is hard to justify until the final answer is found. Therefore, we only verify the [Pertinence] constraint globally for each reasoning step given the full completed reasoning process. When the intermediate reasoning steps are not helpful but even barrier the final answer, the reasoning likely takes more steps to find the final answer. Based on this prior knowledge, we set the maximum reasoning steps as $r$, so that the reasoning is terminated when an answer is not found given more than $r$ steps. We set $r=5$ by default in our experiments.
3.5 Decoding with Depth-first Search
As we have introduced in Section 3.1, there may exist multiple step candidates that satisfy all the constraints. In this case, how can we decide which step is the next step?
To answer this question, let’s first recall how LLMs how to decide which is next token. The most simple method is the greedy decoding, which just takes the highest logit token candiate as the next token. In our case, since we do not have scores to rank step candidates but only the discrete constraint verifications to filter the candidates, we cannot simply take the “best” step as the next reasoning step. Can we have other alternate solutions?
Let’s see a more advanced token-level decoding method: beam search. Beam search takes multiple candidate tokens as the next tokens and store them in a stack to do a best-first search in the token sequence generation. We borrow the idea of the beam search to process the multiple valid step candidates with the depth-first search. Specifically, we store all the step candidates that satisfy the constraints in a stack, as shown in Figure 5. Then, in every iteration, we pop the reasoning step stored at the head of the stack as the next reasoning step and do further reasoning based on it. We name our method as DeepEdit: depth-first search based decoding for knowledge editing. We visualize our DeepEdit in Figure 1(c). Also, we present the iteration-wise outputs of our DeepEdit on an instance in Figure 6.
The strategy of depth-first search has been utilized by some research on LLMs, such as tree-of-thoughts Yao et al. (2023). Different from them, our DeepEdit’s depth-first search is to enforce the LLMs’ decoding to satisfy the constraints derived for KE from a global perspective.
4 New Benchmarks for More Precise and Challenging Evaluation

### v3

@@ Abstract
Answering multi-hop questions involving new knowledge is a challenging task in evaluating large language models’ (LLMs) knowledge editing (KE) methods. This task is rather difficult because the LLMs’ hallucinations on new knowledge would harm the logical coherence of LLMs’ multi-hop reasoning and lead to incorrect answers. To address this issue, we design decoding constraints to “regulate” LLMs’ reasoning, enhancing logical coherence when incorporating new knowledge. We incorporate the constraints into a new KE framework: DeepEdit ( Depth-first Search-based Constrained Decoding for Knowledge Editing), which enhances LLMs to generate coherent reasoning chains with new knowledge through a depth-first search. Our search selects the most important knowledge that satisfies our constraints as the reasoning step to efficiently increase the reasoning depth. In addition to DeepEdit, we propose two new KE benchmarks: MQuAKE-2002 and MQuAKE-hard, which provide more precise and challenging assessments of KE approaches. Qualitatively, DeepEdit enables LLMs to produce succinct and coherent reasoning chains involving new knowledge. Quantitatively, it yields significant improvements on multiple KE benchmarks.
@@ 1 Introduction
Knowledge editing (KE) aims to enable large language models (LLMs) to produce outputs in accordance with the knowledge that is new or conflicted with the knowledge memorized by the LLMs. A challenging task to evaluate KE is the multi-hop question answering with new knowledge Zhong et al. (2023). In this task, LLMs should answer the questions corresponding to a chain of facts including new knowledge (see Figure 1(a)).
Figure 1: (a) An instance of multi-hop question answering with new knowledge. (b) Prior KE methods, e.g., MeLLo Zhong et al. (2023), suffers from the hallucinations of LLM generators. (c) Our DeepEdit controls LLMs with decoding constraints to produce coherent reasoning chains involving new knowledge.
Figure 2: Our DeepEdit controls LLMs’ reasoning to follow the constraints of Conciseness, Coherence, Receptiveness, and Pertinence so as to soundly incorporate new knowledge into LLMs’ question answering.
Recent literature has mainly sought to perform KE through in-context editing Zhong et al. (2023); Cohen et al. (2023); Zheng et al. (2023), which puts new knowledge in the input context to fulfill KE. However, the input-conflicting hallucinations on the in-context new knowledge would harm the logical coherence of multi-step reasoning and lead to incorrect answersZhang et al. (2023).
Taking a closer look at the ground-truth output in Figure 1(a), we notice that: the adjacent reasoning steps should be logically coherent, and new knowledge should participate in the reasoning chain at appropriate positions. This raises a natural question:
@@
The central idea of this paper is to view KE as a problem of constrained decoding. We design the decoding constraints, by following which LLMs can soundly incorporate new knowledge in LLMs reasoning. Specifically, we propose the following constraints: Conciseness requires every reasoning step to be identical; Coherence requires the adjacent reasoning steps to be coherent; Receptiveness requires new knowledge to replace the conflicted memorized knowledge; Pertinence requires every reasoning step to be relevant to the target question (see Figure 2). A multi-step reasoning chain that leads to the correct answer should satisfy the above constraints.
We incorporate the above constraints into a new KE framework: DeepEdit ( Depth-first Search-based Constrained Decoding for Knowledge Editing). When producing every reasoning step, DeepEdit searches over the parametric and new knowledge to find the ones that satisfy our constraints, termed as valid step candidates. Among the valid candidates, DeepEdit takes the most important one as the new reasoning step for further reasoning, which is the (new) knowledge that is most helpful to answer the target question. Our depth-first search strategy efficiently increases the reasoning depth and enhances LLMs to produce the coherent reasoning chains involving new knowledge.
The advantages of our DeepEdit on KE are three-folds. First, it does not need to retrain the LLMs like prior model editing methods Zhu et al. (2020), which is complex and expensive. Our DeepEdit works during inference with negligible costs. Second, our DeepEdit is flexibly applicable to any black-box LLM without requiring access to model parameters or token-wise distributions. Third, our DeepEdit enhances LLMs to soundly utilize new knowledge to produce coherent reasoning chains in accordance with new knowledge.
To provide more precise and challenging assessments of KE methods, we build two new benchmarks MQuAKE-2002, and MQuAKE-hard, which resolve the knowledge-conflicting annotation mistakes in the popular KE benchmark MQuAKE-3k Zhong et al. (2023). We evaluate DeepEdit on the KE benchmarks MQuAKE-3k, MQuAKE-2002, and MQuAKE-hard Zhong et al. (2023); Meng et al. (2022a). Qualitatively, DeepEdit enhances LLMs to produce more succinct and faithful reasoning with new knowledge, as shown in Figure 9. Quantitatively, DeepEdit improves the KE for popular LLMs by a significant margin OpenAI (2022); Touvron et al. (2023).
@@ 2 Related Work
Knowledge Editing
Recent work observes the superior KE performance of in-context editing. MeLLo Zhong et al. (2023) designs a single prompt to alternately conduct the text generation and new knowledge analysis using the LLM generator. Cohen et al. (2023) propose to append the new knowledge the head of the input prompt. In-context editing lacks direct control of the LLMs’ outputs and how the new knowledge influences the outputs is not traceable. Different from them, we propose the new decoding method DeepEdit for KE, which effectively augments LLMs on KE during the decoding stage. We illustrate a comparison between DeepEdit and existing KE methods in Table 1.
Figure 3: When producing every reasoning step, we retrieve the edited facts that are close to the generated step in the semantic space (left). Only the edited facts that satisfy all the constraints, i.e., valid edited facts, are considered as the step candidates (right).
KE Method Editing Editing Editing Model Parameters Input Prompts Outputs MEND Mitchell et al. (2021) MEMIT Meng et al. (2022c) ICE Cohen et al. (2023) MeLLo Zhong et al. (2023) DeepEdit (Ours)
Table 1: Comparison between KE methods.
@@ 3 Methodology
Multi-hop question answering with new knowledge is a challenging task for assessing knowledge editing (KE) of large language models (LLMs) Zhong et al. (2023). This task considers the questions that correspond to a chain of facts including new knowledge. It assumes a database that stores all new facts, which LLMs should adhere to, and posits that only a few facts can influence the ground-truth answer to each question. In this section, we will introduce how we integrate the decoding constraints into the design of a novel KE method, DeepEdit (Depth-first Search based Decoding for Knowledge Editing), to enhance the KE of black-box LLMs.
3.1 Constraint Designing for Knowledge Editing
On a LLM’s multi-step reasoning chain to answer a multi-hop question involving new knowledge, we find the following constraints on reasoning steps can help LLMs to give the correct answer:
@@ 3.1 Constraint Designing for Knowledge Editing
    Conciseness: Every step is identical.
@@
    Coherence: Adjacent steps are coherent.
@@
    Receptiveness: No conflicts exist between reasoning steps and new knowledge.
@@
    Pertinence: Every reasoning step should be relevant to the target question.
Among the above four constraints, Conciseness removes loops in the reasoning chain; Coherence guarantees the coherence of reasoning; Receptiveness ensures the LLMs’ receptiveness to new knowledge; Pertinence guides LLMs to answer the target question. We visualize different constraints in Figure 2. Intuitively, the above constraints are easy to follow for LLMs when there is no new knowledge. However, the LLMs’ utilization of new knowledge is affected by the hallucination, which makes the above constraints easy to break when incorporating new knowledge. In this work, we apply the above constraints to incorporate the new knowledge into LLMs’ decoding.
3.2 Constraint Verification on Every Step
We term the reasoning steps that satisfy all the constraints as valid steps. Because an invalid reasoning step could mislead LLMs into generating incorrect subsequent steps, we verify the constraints during the decoding of each reasoning step.
At every iteration, we take both the parametric and new knowledge as the step candidates. First, we let LLMs generate one reasoning step with temperature as 00, which represents the parametric knowledge that the LLM believes to be useful to answer the target question. Second, we retrieve $N$ new knowledge closest to the generated reasoning step. The distance between knowledge is measured by the semantic distance of sentence-level embeddings given by a pre-trained BERT model Devlin et al. (2019). The closer the context of the new knowledge is to the generated step, the more likely it is that the new knowledge will be helpful in answering the target question. We take these $N+1$ knowledge as the reasoning step candidates, with examples shown in Figure 3.
We verify the constraints on the above step candidates to find the valid step candidates. In principle, every constraint’s verification can be seen as a binary classification problem on a pair of sentences. For example, Coherence’s verification can be completed by assessing whether a pair of reasoning steps is relevant to each other or not. Therefore, one option to do constraint verification is to train a binary classifier for every constraint. However, such a solution is not data or time-efficient. In this work, we develop four specialized in-context learning-based LLM agents to act as the verifiers. Every verifier’s prompt includes $D$ positive and negative demonstrations randomly sampled from the instances of the MQuAKE dataset Zhong et al. (2023). In our experiments, we set $D=2$ and $N=2$ by default.
3.3 Efficient Reasoning with Depth-first Search
At every iteration, there may exist more than valid step candidates. We can take every one of them as the new reasoning step in parallel and make further reasoning based on each of them separately. This design can be seen as a breadth-first search of the LLMs’ multi-hop reasoning chain incorporating new knowledge. It is undoubtedly time-consuming because the time complexity grows exponentially with the number of reasoning steps. This raises a natural question:
Can we more efficiently expand the depth of reasoning chains during decoding?
Our answer is yes. We propose to take the most important step candidate as the reasoning step for further reasoning. Actually, different valid step candidates are not equally important. First, the edited facts should be more important than the generated step, since the former generally contains new knowledge unknown by LLMs. Second, the edited facts that are closer to the generated step are more important, which are more helpful in answering the target question due to their relevance to the LLM’s generated reasoning context. Based on the above prior, in every iteration, we select the most important knowledge as the next reasoning step and do further reasoning based on it, as shown in Figure 4.
Figure 4: At every iteration, we verify the decoding constraints over parametric and new knowledge to find the valid step candidates. Then we take the most important candidate as the next reasoning step to efficiently increase the reasoning depth.
We name our method as DeepEdit: depth-first search-based constrained decoding for knowledge editing. DeepEdit takes the most important step candidate as the new reasoning step at every iteration and efficiently produces the multi-hop reasoning incorporating new knowledge. We summarize DeepEdit in Algorithm 1 and visualize it in Figure 1. Furthermore, we present the iteration-wise outputs of our DeepEdit in Figure 8.
Assume the number of reasoning hops as $h$, and the number of valid step candidates per iteration as $c$. The time complexity of breadth-first search is $c^{h}$. In contrast, our DeepEdit has a much lower time complexity of $\mathcal{O}(h)$ thanks to our depth-first search design considering the importance of different candidates. Empirical results validate DeepEdit’s superior efficiency compared with the breadth-first search.
The strategy of depth-first search has been utilized by some research on LLMs, such as tree-of-thoughts Yao et al. (2023). Different from them, our DeepEdit’s depth-first search is designed to efficiently find the multi-hop reasoning chain incorporating new knowledge. Our DeepEdit is free to be applied to all black-box LLMs: it bypasses the requirement of accessing the token-wise distributions because all the decoding actions are defined at the step level.
Algorithm 1 DeepEdit: Depth-first Search based Constrained Decoding for Knowledge Editing.
Input: The query question $q$. The prompt $p$. The set of new knowledge $\mathcal{K}$. The number of retrieved new knowledge $N$. Maximum number of reasoning depth $r$.
Output: The answer $a$ to the query question.
1:for $i\leftarrow 1$ to $r$ do
2:     $o\leftarrow\mbox{LLM}(p)$
3:     if $o$ is the answer then
4:         return $o$
5:     $\hat{\mathcal{K}}\leftarrow\mbox{search}(o,\mathcal{K},n)$triangleright retrieve the $n$ semantically similar new knowledge to $o$, ranked by the distance to $o$ in the ascending order.
6:     for $j\leftarrow 1$ to $n$ do
7:         $k\leftarrow\hat{\mathcal{K}}[i]$
8:         $nk\leftarrow\mbox{False}$triangleright $nk$ is the boolean variable to decide whether some new knowledge is valid.
9:         if ($k$ satisfies all the constraints) then
10:              $nk\leftarrow\mbox{True}$
11:              $p\leftarrow p+k$
12:              break
13:     if not $nk$ then
14:         $p\leftarrow p+o$      return $o$
4 New Benchmarks for More Precise and Challenging Evaluation of KE